# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v3

@@ I Introduction
In computer vision, deep learning, which makes use of deep convolutional neural networks and their variants, can efficiently extract and accurately classify image features, outperforming traditional methods in tasks such as object detection and image segmentation [5, 6, 7]. In the field of natural language processing, language models that are built on deep learning can automatically learn language grammar, semantics, and contextual associations, performing exceptionally well in tasks such as machine translation and text generation, and propelling the technology towards intelligence and automation [4, 8, 9]. In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11, 12].
@@ II-B Related work
Image reasoning models such as CoPINet[23], LEN+teacher[22], and DCNet[24] primarily focus on learning differences and rules. Meanwhile, models like NCD[26], SAVIR-T[27], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[28, 29, 30]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy. RS-CNN and RS-TRAN[31] demonstrate exceptional capabilities in addressing problems related to RPM, while Triple-CFN[33] stands out among numerous models by implicitly extracting and indexing information about concepts and reasoning units, effectively enhancing reasoning accuracy.
@@ III Methodology
This paper first proposes that to enhance the ability of deep learning models to solve RPM, which are causal reasoning problems, when designing the model’s structure and dividing its modules, one should fully refer to the causal chains within the problems. Based on this methodology, this paper designs a new model for solving RPM problems, named DIO. However, this paper finds that DIO is in satin without substance. Its learning approach is essentially the same as that of the previous RPM solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in RPM instance. This learning approach leads to a situation where, no matter how well its structure aligns with the causal chains in the problems, its learning and optimization will not progress towards the human predefined reasoning framework. Therefore, this paper proposes three methods, namely Brando, WORLD, and DIEGO, to refine DIO’s learning objectives. As a bonus, when DIO adopts the WORLD method, it demonstrates the capability to solve generative RPM problems.
@@ IV DIO: A Baseline Model
In this section, we design a high-performance RPM solving model, named DIO (Deep-learning Intelligent-model Organized via Causal Chains). While providing a detailed exposition of the model’s structure, we have also elucidated its underlying design principles.
@@
This paper proposes that designing the structure of an RPM solving model in accordance with this causal chain will yield high performance, which prompted us to design the DIO model. Accordingly, the DIO comprises four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
@@ IV-D The Progressive Pattern Consistency Evaluation Module
V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM Solving Model
@@ V-A The learning objective of the DIO
The right-hand side of Inequality (14) is precisely the classical formulation for optimizing the probabilistic distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the Cross-Entropy (CE) loss function. From this, it can be inferred that for those past end-to-end RPM solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
@@
In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options. When optimizing the DIO using $\ell_{\text{Brando}}$ as the loss function, it can induce bounded hypothetical options while leveraging a tighter variational lower bound to enable DIO to maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
@@ VI-B Map Out Hypothetical Options
Formulas (18) and (19) indicate that hypothetical options should demonstrate constructiveness for any given context matrix, meaning that hypothetical options should be customized according to the given context matrix. Therefore, training a set of shared hypothetical options is inappropriate. This implies that we require a mapping with the following form:
@@ VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
VIII DIEGO: Methods for Rectifying the Causal Chain in DIO
@@
VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
This paper realizes that setting maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. Although this mutual information measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic, yielding a model optimized for high accuracy rather than for faithful causal reasoning. Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet neither resolves this limitation. In a nutshell, mutual-information-driven approaches remain effective but must be complemented by an explicit causal perspective.
Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled. The network architecture of DIO is designed to mirror this causal chain, aiming for DIO to automatically capture the chain during training. Nevertheless, while this design guarantees structural alignment, it cannot ensure that the latent variables derived by DIO are also semantically aligned. Given that mutual information fails to account for causality, the semantic alignment awaits empirical verification.
In order to verify semantic alignment, this paper selects the “Distribute 9” dataset from the RPM-style RAVEN benchmark [19] as the research object. We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD. Subsequently, we select 30 instances with strictly identical progressive patterns from this dataset and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances. The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
@@ VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
If DIO accurately models the causal chain embedded in RAVEN problems, then once DIO encodes 30 instances sharing the same progressive patterns into pattern representations, those representations should exhibit high cosine similarity among themselves. Contrary to expectations, the three similarity matrices in Figure 9 fail to support this anticipated outcome, indicating that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO systematically deviates from the ground-truth causal chains inherent in RPM problems. This underscores the limitations of adopting mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective. Given these findings, we propose that further modeling of the causal chain could start with semantic rectification in DIO, which would substantially enhance its performance in solving RPM problems.
@@
It is widely accepted that introducing more supervisory signals should improve model performance. Indeed, previous studies have attempted to leverage metadata as additional supervision by designing progressive pattern-learning tasks, aiming to boost both interpretability and inference accuracy simultaneously. Recent work [25, 28, 27, 31] has collectively shown, however, that this practice can backfire: directly embedding metadata into proposed models actually degrades reasoning performance. This paper argues that the DIO model, constructed upon causal chains, together with the DIEGO method that strategically rectifies key segments of those chains, offers a viable way to overcome this limitation.
@@ IX-A WORLD Guides DIO on Generative RPM problems
In essence, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
@@ IX-B Evaluating the Generated Solutions
More importantly, WORLD is a strong mutual-information-driven inductive model. Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the rationality and potential of mutual information as a reasoning cue.
@@ X-1 Solo DIO
which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies. In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
@@ X-2 Brando-enhanced DIO
which maps out constructive hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-16=8$) constructive hypothetical incorrect options to the DIO in each iteration;
@@ X-3 WORLD-augmented DIO
utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tightening the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the WORLD framework, the hyperparameter $M$ is set to 256, equipping the GMM employed for DIO with 256 components. This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options;
@@ X-4 DIEGO-guided DIO
incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns. In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem.
@@ X-B Experiment on PGM Problem
TABLE III: Reasoning Accuracies on PGM Generalization Tasks.
@@ XI-B Ablation Study on Brando
The ablation study on Brando focuses on the hyperparameter $\beta$. $\beta$ denotes the number ($\beta-17$) of hypothetical options the Brando network provides to DIO. As this paper shows that the count of constructive incorrect options for DIO directly influences the tightness of its derived mutual information variational lower bound, which in turn directly impacts DIO’s reasoning potential. Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VII.
TABLE VII: Reasoning Accuracies of DIO+Brando under Different $\beta$ Values.
@@
Table VII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved. The asterisk (*) on a sub-task indicates that, when the Brando network is applied to this task, the process of learning historical image features defined by Equation (27) is deactivated, facilitating ablation and analysis of this mechanism. The asterisk-marked results demonstrate that learning historical image features is effective in mapping out constructive hypothetical options.
@@ XI-C Ablation Study on WORLD
The ablation study on WORLD focuses on the hyperparameter $M$. $M$ indicates the number of components in the GMM constructed by World for RPM image features. The value of $M$ determines the maximum diversity of solution distributions that WORLD can generate for DIO, expressed as $M^{N}$. Therefore, ablation experiments centered around the value of $M$ are warranted. The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table VIII.
TABLE VIII: Reasoning Accuracies of DIO+WORLD under Different $M$ Values.
@@
PGM Neutral 99.2 99.0 98.5 98.0 98.1
The number of components directly affects the complexity of the GMM. WORLD proposes an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM. As a highly versatile distribution model, the GMM can theoretically ensure WORLD’s broad applicability. However, as shown in Table VIII, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task. This paper contends that this phenomenon arises because a 32-component GMM fails to capture the complexity of the underlying distribution of PGM image features. If DIO attempts to fit these more complex features into such a simplified GMM, a performance decline can be anticipated. Therefore, this study establishes a critical prerequisite for WORLD’s applicability: the GMM modeled by WORLD must be able to adapt to the inherent complexity of the target problem.
In traditional deep learning tasks, underfitting is a prevalent phenomenon arising when neural networks lack sufficient capacity. However, WORLD makes this phenomenon manageable through its $M$-dependent capacity control. Therefore, particular attention is required to the parameter $M$ when applying this method.

### v4

@@ I Introduction
In computer vision, deep learning, which makes use of deep convolutional neural networks and their variants, can efficiently extract and accurately classify image features, outperforming traditional methods in tasks such as object detection and image segmentation [5, 6, 7]. In the field of natural language processing, language models that are built on deep learning can automatically learn language grammar, semantics, and contextual associations, performing exceptionally well in tasks such as machine translation and text generation, and propelling the technology towards intelligence and automation [4, 8, 9]. In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11].
@@ II-B Related work
Image reasoning models such as CoPINet[21] and DCNet[22] primarily focus on learning differences and rules. Meanwhile, models like NCD[24], SAVIR-T[25], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[26, 27, 28]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy. RS-CNN and RS-TRAN[29] demonstrate exceptional capabilities in addressing problems related to RPM, while Triple-CFN[31] stands out among numerous models by implicitly extracting and indexing information about concepts and reasoning units, effectively enhancing reasoning accuracy.
@@ III Methodology
This paper first proposes that to enhance the ability of deep learning models to solve RPM, which are causal reasoning problems, when designing the model’s structure and dividing its modules, one should fully refer to the causal chains within the problems. Based on this methodology, this paper designs a new model for solving RPM problems, named DIO. However, this paper finds that DIO is in satin without substance. Its learning approach is essentially the same as that of the previous RPM-solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in the RPM instance. This learning approach leads to a situation where, no matter how well its structure aligns with the causal chains in the problems, its learning and optimization will not progress towards the human predefined reasoning framework. Therefore, this paper proposes three methods, namely Brando, WORLD, and DIEGO, to refine DIO’s learning objectives. As a bonus, when DIO adopts the WORLD method, it demonstrates the capability to solve generative RPM problems.
@@ IV DIO: A Baseline Model
In this section, we design a high-performance RPM-solving model, named DIO (Deep-learning Intelligent-model Organized via Causal Chains). While providing a detailed exposition of the model’s structure, we have also elucidated its underlying design principles.
@@
This paper proposes that designing the structure of an RPM-solving model in accordance with this causal chain will yield high performance, which prompted us to design the DIO model. Accordingly, the DIO comprises four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
@@ IV-D The Progressive Pattern Consistency Evaluation Module
V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM-solving Model
@@ V-A The learning objective of the DIO
The right-hand side of Inequality (14) is precisely the classical formulation for optimizing the probabilistic distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the Cross-Entropy (CE) loss function. From this, it can be inferred that for those past end-to-end RPM-solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM-solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
@@
In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive ones. By adopting $\ell_{\text{Brando}}$ as the training loss function, DIO not only obtains constructive hypothetical options but also parameterizes a tighter variational lower bound, thereby maximizing the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
@@ VI-B Map Out Hypothetical Options
Formulas (18) and (19) indicate that hypothetical options should demonstrate constructiveness for any given context matrix, meaning that such options should be customized according to the provided context matrix. This implies that naively training a set of shared hypothetical options across all context matrices is suboptimal. Given this, we require a mapping with the following form:
@@ VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
VIII DIEGO: Methods for Rectifying the Causal Chain modeled by DIO
@@
VIII-A Target Reasoning Rather Than Solving Problems
This paper realizes that setting maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. Although this statistical measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic [35], yielding a model optimized for high accuracy rather than for faithful causal reasoning. Thus, the performance gains brought by mutual information may be superficial. Both the Brando and WORLD methodologies aim to assist DIO in maximizing mutual information, yet neither resolves this limitation. In a nutshell, mutual-information-driven approaches are effective but need an explicit causal perspective.
Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled. The network architecture of DIO is designed to mirror this causal chain, aiming for DIO to automatically capture the chain during training. Nevertheless, while this design guarantees structural alignment, it cannot ensure that the latent variables derived by DIO are also semantically aligned. Given that mutual information fails to account for causality [35], the semantic alignment awaits empirical verification.
In order to verify semantic alignment, this paper takes the “Distribute 9” dataset from the RPM-style RAVEN benchmark [18] as the research object. We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD. Subsequently, we extract 30 instances that share strictly identical progressive patterns and compute the pairwise similarities among the progressive pattern representations that DIO encodes for these 30 instances. The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
@@ VIII-A Target Reasoning Rather Than Solving Problems
If DIO accurately models the causal chain embedded in RAVEN problems, then once DIO encodes 30 instances sharing the same progressive patterns into pattern representations, those representations should exhibit high cosine similarity among themselves. Contrary to expectations, the three similarity matrices in Figure 9 fail to support this anticipated outcome. This finding indicates that, even with auxiliary supervision from Brando or WORLD, the causal chain learned by DIO systematically diverges from the ground-truth chain inherent in RPM problems. This further underscores the limitations of adopting mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective. Therefore, we propose that further modeling of the causal chain could prioritize the rectification of DIO’s semantic representations, which would substantially enhance its performance on RPM problems.
@@
It is widely accepted that introducing more supervisory signals should improve model performance. Indeed, previous studies have attempted to leverage metadata as additional supervision, aiming to boost both interpretability and reasoning accuracy simultaneously. Recent work [23, 26, 25, 29] has collectively shown, however, that this practice can backfire: directly embedding metadata into proposed models actually degrades reasoning performance. This paper argues that the DIO model, constructed upon causal chains, together with the DIEGO method that strategically rectifies key segments of those chains, offers a viable way to overcome this limitation.
@@ IX-A WORLD Guides DIO on Generative RPM problems
In a nutshell, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
@@ IX-B Evaluating the Generated Solutions
More importantly, WORLD is a strong mutual-information-driven inductive model, yet it cannot escape the inherent drawback of neglecting reasoning logic. Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the extent to which WORLD’s performance gains are merely superficial.
@@ X-1 Solo DIO
In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
@@ X-2 Brando-enhanced DIO
In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-16=8$) constructive hypothetical incorrect options to the DIO in each iteration;
@@ X-3 WORLD-augmented DIO
In the WORLD framework, the hyperparameter $M$ is set to 512, equipping the GMM employed for DIO with 512 components. This configuration enables the generation of $M^{N}=512^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options. It is worth noting that when WORLD is used to assist DIO in solving generative RPM problems, $M$ is set to 2048;
@@ X-4 DIEGO-guided DIO
In the DIEGO framework, the parameter $F$ is determined by the number of metadata modalities in the targeted RPM problem.
@@ X-B Experiment on PGM Problem
TABLE III: Accuracies on PGM Generalization Tasks.
@@ XI-A Ablation Study on DIO
Moreover, across all RAVEN and PGM sub-tasks, DIO uniformly splits every image into $4\times 4$ patches. This design deliberately bypasses the grid-like entity layouts in RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks, preventing unintended inductive biases and prior injection while embracing the causal chain rationale of RPM problems: multiple entities must be analyzed holistically. We therefore conducted ablation experiments on the hyperparameter $N$ to examine how aligning the patch grid with RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks affects performance, and the results are shown in Table VII. The results indicate that favoring priors over the causal chain may undermine model performance.
TABLE VII: The reasoning accuracy of DIO when its patch grid is aligned with entity positions in the “2×2 Grid” and “3×3 Grid” sub-tasks of RAVEN and I-RAVEN.
Accuracy(%)
Patch grid 2 $\times$ 2 Grid 3 $\times$ 3 Grid
Aligned with entity 95.0/98.3 92.8/94.2
$4\times 4$ Grid 99.6/99.7 96.1/97.4
@@ XI-B Ablation Study on Brando
The ablation study on Brando focuses on the hyperparameter $\beta$. $\beta$ denotes the number ($\beta-17$) of hypothetical options the Brando network provides to DIO. As this paper shows that the count of constructive incorrect options for DIO directly influences the tightness of its derived mutual information variational lower bound, which in turn directly impacts DIO’s reasoning potential. Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VIII.
TABLE VIII: Accuracies of DIO+Brando under Different $\beta$.
@@
Table VIII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved. The asterisk (*) on a sub-task indicates that, when the Brando network is applied to this task, the process of learning historical image features defined by Equation (27) is deactivated, facilitating ablation and analysis of this mechanism. The asterisk-marked results demonstrate that learning historical image features is effective in mapping out constructive hypothetical options.
@@ XI-C Ablation Study on WORLD
The ablation study on WORLD focuses on the hyperparameter $M$. $M$ indicates the number of components in the GMM constructed by World for RPM image features. The value of $M$ determines the maximum diversity of solution distributions that WORLD can generate for DIO, expressed as $M^{N}$. Therefore, ablation experiments centered around the value of $M$ are warranted. The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table IX.
TABLE IX: Accuracies of DIO+WORLD under Different $M$.
@@
PGM Neutral 99.3 99.0 98.5 98.0 98.1
The number of components directly affects the complexity of the GMM. WORLD proposes an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM. As a highly versatile distribution model, the GMM can theoretically ensure WORLD’s broad applicability. However, as shown in Table IX, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task. This paper contends that this phenomenon arises because a 32-component GMM fails to capture the complexity of the underlying distribution of PGM image features. If DIO attempts to fit these more complex features into such a simplified GMM, a performance decline can be anticipated. Therefore, this study establishes a critical prerequisite for WORLD’s applicability: the GMM modeled by WORLD must be able to adapt to the inherent complexity of the target problem.